Find functions f and g so that h(x)=f(g(x)). The function g must be a polynomial or rational function.
h(x)= ln(2x+5)
Thanks in advance to anyone who helps, I really appreciate it!

Answer by jsmallt9(3758)

Some possible answers:
f(x) = ln(x) g(x) = 2x+5
f(x) = ln(x+5) g(x) = 2x
f(x) = ln((1/2)x + 5) g(x) = 4x
f(x) = ln(x+10) g(x) = 2x-5
f(x) = ln((2/x) + 5) g(x) = 1/x
f(x) = ln((1/x) + 5) g(x) = 1/(2x)
etc.
All of these work and all of them fit the requirement that g be a polynomial or rational function. (The first one is probably the one your teacher is looking for.)